# Training objective drives the consistency of representational similarity across datasets

### v1

@@ Appendix B Downstream task performance
We evaluated each model’s downstream task performance on every dataset by calculating the top-1 accuracy of a linear probe trained on features extracted from the corresponding training set. Each model was evaluated with 3 different random seeds on each dataset, and the mean top-1 accuracy across seeds was used. Hyperparameter selection was performed with a validation set of $20\%$ of the training set and the remaining training data for optimization. We followed the binary search procedure described in [65] and searched for the optimal weight decay parameter $\lambda$ in the interval between $10^{-6}$ and $10^{2}$ in 96 logarithmically spaced steps. This was done for all learning rates $\eta\in\{10^{i}\}_{i=1}^{4}$. After hyperparameter selection, the linear probe was retrained on the full training set and evaluated on the respective test set. All linear probes are trained for 20 epochs, using the AdamW optimizer [54] and a cosine schedule for learning rate decay [53].
Appendix C CKA sensitivity to number of samples in dataset
@@
Our analysis of model similarities requires reliable CKA measurements while managing computational constraints. To identify the minimum number of samples needed for stable CKA values, we analyzed ImageNet-1k subsets with varying samples per class ($k\in{1,5,10,20,30,40}$). We generated pairwise similarity matrices (Fig. 9) and assessed stability by comparing the absolute differences between consecutive sample sizes (Fig. 10).
The results show that most CKA variants converged with 10,000 samples (k=10), with only CKA RBF ($\sigma=0.2$) requiring more samples for stability. Based on these findings, we set k=10 as our standard sample size and applied stratified subsampling to limit datasets larger than 10,000 samples. For all analyses, we selected CKA RBF ($\sigma=0.4$) and CKA linear as our local and global measures, respectively. While this evaluation used a slightly different set of models than our main analysis, it provides guidance about the number of samples that ought to be used to guarantee the stability of CKA measurements.
@@
Figure 9: Pairwise model representational similarity matrices using different dataset sizes (1-40 samples per ImageNet-1k class) and metrics (CKA RBF 0.2-0.8 to linear). Rows indicate increasing samples per class, while columns progress from local to global similarity measures
@@
Figure 10: Distribution of absolute differences of pairwise model representational similarities for each CKA variant measured on datasets with consecutive sample sizes per class ($1{\rightarrow}5$, $\cdots$, $30{\rightarrow}40$). Most CKA variants converge at $k=10$ samples per class, with CKA RBF ($\sigma=0.2$) requiring larger sample sizes for stability.
Appendix D Variation of CKA values near the upper bound
@@
The CKA similarity measure is bounded between 0 and 1, with values close to 1 indicating highly similar representations. However, this upper bound might lead to saturation effects. To better differentiate between highly similar model representations, we apply two nonlinear transformations: $\arccos$ and $\tan$. Both transformations expand the range of high CKA values, with $\tan$ slightly reducing variation for low similarity values. These transformations amplify the inverse U-shape relationship noted in Section 4.2 (Figure 11), suggesting that the relationship between mean and variance of similarities is robust across different scales.
@@
Appendix E Does local or global similarity differ?
@@
As shown in the main text (Fig. 3), mean representational similarities across datasets demonstrate strong positive correlations when measured with different similarity metrics. In this section, we verify that our findings about similarity consistency remain valid when using both CKA RBF 0.4 and RSA Spearman metrics rather than just CKA linear.
To validate this stability across metrics, we correlate the consistency values across all dataset pairs and model set pairs. As shown in Fig. 12, there is a strong positive correlation between global (CKA linear) and local (CKA RBF with $\sigma=0.4$) measures, indicating that similarity consistency observations based on CKA linear can be generalized to the local measure. While the correlation between the two global measures (CKA linear and RSA Spearman) is somewhat lower, it still shows moderate correlation. Even with this reduced correlation strength, similarity consistency measured using RSA Spearman exhibits the same patterns as CKA linear, particularly regarding the model category we found to be most influential – the training objective (see Fig. 13). This additional analysis suggests that our observations on similarity consistency do not depend on the metric used.
@@
Figure 12: Similarity consistency values measured with different metrics. Left: CKA RBF 0.4 versus CKA linear. Right: RSA Spearman versus CKA linear. Correlations are measured using Pearson correlation. Each point represents the consistency value for a pair of datasets and model sets.
@@
Appendix F Validating training objective effects through controlled anchor model analysis
@@
To validate our findings regarding the influence of different model categories (Sec. 4.5), we conducted a controlled analysis using six carefully selected anchor models: OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, and ViT-L. These models were chosen to create systematic variations across training objectives (image-text, self-supervised, and supervised learning), while controlling for architecture (convolutional ResNet-50 vs transformer ViT-L) and model size. This controlled setup helps verify whether the patterns observed in our main analysis persist when comparing individual representative models to broader model sets. For each anchor model, we define a single-element model set (e.g., $\Phi_{\text{OpenCLIP RN50}}$) and compute its similarity consistency with different model groups (e.g., ${\bm{R}}(\Phi_{\text{OpenCLIP RN50,}},{\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}})$.
@@
Fig. 14 shows the correlation coefficient distributions across dataset pairs, focusing on training objectives as they emerged as the most influential factor for similarity consistency. The results support our main findings while providing additional insights. Self-supervised models (SimCLR RN50, DINOv2 ViT-L) consistently show the highest median correlations and lowest variations when compared with any anchor model, regardless of the anchor’s training objective. Surprisingly, even image-text (OpenCLIP RN50, OpenCLIP ViT-L) and supervised (ResNet-50, ViT-L) anchor models correlate more strongly with ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$than with their own categories (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$and ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$, respectively). We also observe a notable pattern of weak correlations between image-text anchors and supervised models (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) and vice versa. These patterns remain remarkably consistent across different architectures and model sizes, suggesting that training objectives, rather than architectural choices or model sizes, primarily drive similarity relationships across datasets.
@@
Figure 14: Distribution of Pearson correlation coefficients between anchor models (OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, ViT-L) and models trained with different objectives. Each box shows correlations across all dataset pairs between one anchor model and models from specific sets (e.g., ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$). Results for CKA RBF ($\sigma=0.4$) and CKA linear are separated by dotted lines. Subplot titles are colored according to the anchor model’s subcategory.

### v2

@@ Appendix B Downstream task performance
We evaluated each model’s downstream task performance on every dataset by calculating the top-1 accuracy of a linear probe trained on features extracted from the corresponding training set. Each model was evaluated with 3 different random seeds on each dataset, and the mean top-1 accuracy across seeds was used. Hyperparameter selection was performed with a validation set of $20\%$ of the training set and the remaining training data for optimization. We followed the binary search procedure described in (Radford et al., 2021) and searched for the optimal weight decay parameter $\lambda$ in the interval between $10^{-6}$ and $10^{2}$ in 96 logarithmically spaced steps. This was done for all learning rates $\eta\in\{10^{i}\}_{i=1}^{4}$. After hyperparameter selection, the linear probe was retrained on the full training set and evaluated on the respective test set (validation set for Imagenet-1k). All linear probes are trained for 20 epochs, using the AdamW optimizer (Loshchilov & Hutter, 2019) and a cosine schedule for learning rate decay (Loshchilov & Hutter, 2017).
Appendix C CKA sensitivity to the number of samples in dataset
@@
Our analysis of model similarities requires reliable CKA measurements while managing computational constraints. To identify the minimum number of samples needed for stable CKA values, we analyzed ImageNet-1k subsets with varying samples per class ($k\in{1,5,10,20,30,40}$). We generated pairwise similarity matrices (Fig. 10) and assessed stability by comparing the absolute differences between consecutive sample sizes (Fig. 11). The results show that most CKA variants converged with $10{,}000$ samples (k=10), with only CKA RBF ($\sigma=0.2$) requiring more samples ($30{,}000$, k=30) for stability.
To further assess CKA stability across different samples, we conducted bootstrapping experiments using six representative models (OpenCLIP ViT-L, OpenCLIP RN50, ResNet-50, ViT-L, DINOv2 ViT-L, and SimCLR RN50—the same anchor models described in Appx. G). For CKA linear, we performed 500 bootstrap iterations with $10{,}000$ images per subset. For CKA RBF ($\sigma=0.2$), we performed 500 bootstrap iterations with $30{,}000$ images per subset. For each bootstrap sample, we extracted features from all six models and computed CKA values between each model pair.
Fig. LABEL:fig:append-bootstrap-cka-linear confirms the stability of CKA linear values across all model pairs, while for CKA RBF ($\sigma=0.2$), Fig. LABEL:fig:append-bootstrap-cka-rbf-02 reveals that even when using the larger $30{,}000$ samples per subset, we still observe slightly higher variability compared to CKA linear.
These findings indicate that local similarity measures (RBF kernel) are more sensitive to specific stimuli than global similarity measures (linear kernel). This aligns with theoretical expectations, as stimulus-specific fine-grained details have a greater influence on local similarity measurements.
Based on these findings, we applied stratified subsampling to limit all datasets to $10{,}000$ or $30{,}000$ samples for CKA linear and CKA RBF ($\sigma=0.2$), respectively.
@@
Figure 10: Pairwise model representational similarity matrices using different dataset sizes (1-40 samples per ImageNet-1k class) and metrics (CKA RBF 0.2-0.8 to linear).
@@
Figure 11: Distribution of absolute differences of pairwise model representational similarities for each CKA variant measured on subsets of ImageNet-1k with consecutive sample sizes per class ($1{\rightarrow}5$, $\cdots$, $30{\rightarrow}40$). Most CKA variants converge at $k=10$ samples per class, with CKA RBF ($\sigma=0.2$) requiring larger sample sizes for stability.
Appendix D Similarity matrices for different datasets
In § 4.2, we showed that similarity measurements are not directly transferable across datasets due to their high variability. Fig.  2 displays representational similarity matrices computed using CKA linear across three selected datasets, along with the mean and standard deviation matrices across all evaluation datasets. In Fig. 13, we show additional similarity matrices for three representative datasets from each dataset category (see Appx. A for details). The yellow-boxed models, containing image-text models, are most similar on natural image datasets. In contrast, the white-boxed models, containing self-supervised models, show the highest similarity on specialized EuroSAT and PCAM datasets. The cyan-boxed models, containing supervised models, are most dataset dependent, showing high similarity on Pets but low similarity on ImageNet-1k.
Figure 13: Representational similarity of model pairs, calculated with linear CKA. For each dataset category, three evaluation datasets are displayed.
Appendix E Variation of CKA values near the upper bound
@@
The CKA similarity measure is bounded between 0 and 1, with values close to 1 indicating highly similar representations. However, this upper bound might lead to saturation effects. To better differentiate between highly similar model representations, we apply two nonlinear transformations: $\arccos$ and $\tan$. Both transformations expand the range of high CKA values, with $\tan$ slightly reducing variation for low similarity values. These transformations amplify the inverse U-shape relationship noted in Section 4.2 (Fig. 14), suggesting that the relationship between mean and variance of similarities is robust across different scales.
@@
Appendix F Influence of similarity metric on similarity consistency
@@
In our main analysis, we used CKA linear as a global similarity metric. Fig. 3 shows strong positive correlations between mean representational similarities across datasets when measured with different metrics. Here, we investigate whether this stability extends to similarity consistency distributions when using alternative metrics.
@@
Figure 15: Relationship between similarity consistency values measured with different similarity metrics across all model sets and dataset pairs. From left to right: CKA linear vs. CKA RBF ($\sigma=0.4$), CKA linear vs. CKA RBF ($\sigma=0.2$), CKA RBF ($\sigma=0.2$) vs. CKA RBF ($\sigma=0.4$), and CKA linear vs. RSA Spearman. Correlations are measured using Pearson correlation. Each point represents the similarity consistency value for a pair of datasets and model sets.
Local similarity metrics
When examining local similarity structures, we leverage the flexibility of CKA with an RBF kernel. The kernel bandwidth parameter $\sigma$ controls sensitivity to local structure, allowing us to capture different aspects of representational similarity. Notably, Fig. 10 reveals that pattern differences between CKA linear and CKA RBF ($\sigma=0.2$) are substantially more pronounced than those between CKA linear and CKA RBF ($\sigma=0.4$). This suggests that the latter partially captures global similarity structures rather than purely local ones. Indeed, the correlation between similarity consistency values measured with CKA linear and CKA RBF ($\sigma=0.4$) is strong (Fig. 15, leftmost panel), while it is substantially lower between CKA linear vs. CKA RBF ($\sigma=0.2$) (Fig. 15, second panel). The correlation between similarity consistency values measured with the two different sigmas ($\sigma=0.2$ and $\sigma=0.4$) (Fig. 15, third panel) is in between, strengthening the hypothesis that CKA RBF ($\sigma=0.4$) partially captures global structure. Therefore, we use CKA RBF ($\sigma=0.2$) as our local similarity metric.
We observe the same overall pattern as in Fig. 6 of the main text for the CKA RBF ($\sigma=0.2$) similarity metric: the objective is a main influencing factor for similarity consistency, while network architecture and model size are less important (Fig. 16). However, the distributions of consistency values reveal two interesting observations. For local similarity, supervised model pairs (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) are more consistent than the image-text model pairs (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$), which are better at representing global structure. In addition, models trained on IN21k $({\color[rgb]{1,0.37,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.37,1}\pgfsys@color@cmyk@stroke{0}{0.63}{0}{0}\pgfsys@color@cmyk@fill{0}{0.63}{0}{0}\Phi_{\text{IN21k}}},{\color[rgb]{1,0.37,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.37,1}\pgfsys@color@cmyk@stroke{0}{0.63}{0}{0}\pgfsys@color@cmyk@fill{0}{0.63}{0}{0}\Phi_{\text{IN21k}}})$ are more consistent than models trained on IN1k $({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}})$. Their ordering flipped compared to Fig. 6, where $({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}})$ model pairs are more consistent in their global structure. IN21k contains substantially more classes (21.843) and a higher percentage of the classes are leaf nodes in the WordNet (Princeton University, 2010) tree (76.71% for IN21k vs. 65% for IN1k), representing more fine-grained entities. The representation must contain more fine-grained details to distinguish these classes, dominating local similarities.
Other global similarity metrics
In Fig. 17, the correlation between the two global measures (CKA linear and RSA Spearman) is moderate. Even with this reduced correlation strength, similarity consistency measured using RSA Spearman exhibits the same patterns as CKA linear, particularly regarding the model category we found to be most influential – the training objective (see Fig. 17).
The analysis presented in this section suggests that our observations on the influence of the training objective on similarity consistency hold for multiple metrics.
Figure 16: Distribution of similarity consistencies for each model set pair based on CKA RBF ($\sigma=0.2$) representational similarities. The boxes are sorted in decreasing median correlation. The dotted line indicates the overall median, while the gray area spans the 25th to 75th percentiles of correlations across all model pairs.
@@
Appendix G Validating training objective effects through controlled anchor model analysis
@@
To validate our findings regarding the influence of different model categories (§4.5), we conducted a controlled analysis using six carefully selected anchor models: OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, and ViT-L. These models were chosen to create systematic variations across training objectives (image-text, self-supervised, and supervised learning), while controlling for architecture (convolutional ResNet-50 vs transformer ViT-L) and model size. This controlled setup helps verify whether the patterns observed in our main analysis persist when comparing individual representative models to broader model sets. For each anchor model, we define a single-element model set (e.g., ${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{OpenCLIP RN50}}}$) and compute its similarity consistency with different model groups (e.g., ${\bm{R}}({\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{OpenCLIP RN50}}},{\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}})$.
@@
Fig. 18 shows the correlation coefficient distributions across dataset pairs, focusing on training objectives as they emerged as the most influential factor for similarity consistency. The results support our main findings while providing additional insights. Self-supervised models (SimCLR RN50, DINOv2 ViT-L) consistently show the highest median correlations and lowest variations when compared with any anchor model, regardless of the anchor’s training objective. Surprisingly, for global similarity, even image-text (OpenCLIP RN50, OpenCLIP ViT-L) and supervised (ResNet-50, ViT-L) anchor models correlate more strongly with ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$than with their own categories (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$and ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$, respectively). However, the image-text anchors show high variation of similarity consistency for local similarity measures and all training objectives. We also observe a notable pattern of weak correlations between image-text anchors and supervised models (${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) and vice versa. These patterns remain remarkably consistent across different architectures and model sizes, suggesting that training objectives, rather than architectural choices or model sizes, primarily drive similarity relationships across datasets.
@@
Figure 18: Distribution of Pearson correlation coefficients between anchor models (OpenCLIP RN50, OpenCLIP ViT-L, SimCLR RN50, DINOv2 ViT-L, ResNet-50, ViT-L) and models trained with different objectives. Each box shows correlations across all dataset pairs between one anchor model and models from specific sets (e.g., ${\color[rgb]{1,0.49,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49,0}\pgfsys@color@cmyk@stroke{0}{0.51}{1}{0}\pgfsys@color@cmyk@fill{0}{0.51}{1}{0}\Phi_{\text{SSL}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$). Results for CKA RBF ($\sigma=0.2$) and CKA linear are separated by dotted lines. Subplot titles are colored according to the anchor model’s subcategory.
Appendix H Effect of evaluation data on similarity consistency
In the main analysis, we examined similarity consistency across model categories for all evaluation datasets (§4.5) and across individual datasets for all models (§4.6). Here, we investigate whether model category effects differ between natural and non-natural (specialized and structured) datasets, as non-natural datasets represent out-of-distribution data for many models (Fig. LABEL:fig:effect-natural-non-natural-data).
The primary finding from §4.5 holds: training objective remains the dominant factor influencing similarity consistency regardless of dataset type. However, similarity consistency between image-text and supervised models (${\color[rgb]{0.0600000000000001,0.46,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.0600000000000001,0.46,1}\pgfsys@color@cmyk@stroke{0.94}{0.54}{0}{0}\pgfsys@color@cmyk@fill{0.94}{0.54}{0}{0}\Phi_{\text{Img-Txt}}}$, ${\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\Phi_{\text{Sup}}}$) shows lower medians and higher variance on natural datasets compared to non-natural ones. For natural datasets, supervised models might encode features tightly coupled to class-label supervision, which diverge from the semantically richer, language-aligned representations of image-text models. Conversely, on non-natural datasets where both model types operate further from their training distributions, this representational gap narrows as both might rely on more generalized features. Complementary, the analysis—examining how training dataset composition affects these patterns—is presented in the next section.
Appendix I Effect of training data domain on similarity consistency
In our main analysis, we restricted our focus to general-purpose vision models trained on large-scale datasets with diverse semantic classes (§ 4.1). This focus was motivated by the Platonic representation hypothesis, which suggests that as the semantic diversity of training data approaches the complexity of reality itself, representations should increasingly converge (Huh et al., 2024). However, Conwell et al. (2024) demonstrated that training data is a key driver of representational similarity between models, potentially more significant than architecture, training objective, or choice of similarity metric. In this section, we examine how this observation manifests in terms of representational consistency.
To investigate the role of training data domain, we selected four architectures (AlexNet (Krizhevsky et al., 2012a), DenseNet161 (Huang et al., 2017), ResNet18 (He et al., 2016), and ResNet50 (He et al., 2016)) with available pre-trained weights for both the general-purpose ImageNet-1k and the scene-specific Places365 (Zhou et al., 2018) datasets. Models were sourced from torchvision [^1] https://pytorch.org/vision/stable/index.html, using default ImageNet weights and Places365 weights from the official GitHub repository [^2] https://github.com/CSAILVision/places365. This selection of models allowed us to isolate the training data effect on similarity consistency while fixing the other characteristics: training objective (Supervised), architecture type (CNN), and model size (small).
We evaluated similarity consistency across the 23 downstream datasets for all model set pairs: $({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}})$, $({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.35,0.87,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.35,0.87,1}\pgfsys@color@cmyk@stroke{0.65}{0.13}{0}{0}\pgfsys@color@cmyk@fill{0.65}{0.13}{0}{0}\Phi_{\text{Places365}}})$, $({\color[rgb]{0.35,0.87,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.35,0.87,1}\pgfsys@color@cmyk@stroke{0.65}{0.13}{0}{0}\pgfsys@color@cmyk@fill{0.65}{0.13}{0}{0}\Phi_{\text{Places365}}},{\color[rgb]{0.35,0.87,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.35,0.87,1}\pgfsys@color@cmyk@stroke{0.65}{0.13}{0}{0}\pgfsys@color@cmyk@fill{0.65}{0.13}{0}{0}\Phi_{\text{Places365}}})$, and $({\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}},{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}})$, where ${\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}}$ contains all models.
Our analysis reveals two main observations. First, as shown in Fig. LABEL:fig:appenddistr_r_in1k_vs_places, Places365-trained model pairs demonstrate more consistent representational similarities compared to ImageNet-1k-trained pairs We hypothesize that domain-specific training constrains the solution space more tightly, whereas the greater diversity of semantic classes in ImageNet-1k allows for more varied learned representations across models. This distinction is particularly interesting in light of the platonic representation hypothesis. While domain-specific models show higher consistency, this may reflect convergence to specialized solutions rather than to general representations of reality. In contrast, the lower consistency among ImageNet-1k-trained models likely reflects the challenge of learning representations that capture broad semantic diversity. This consistency might be even lower for models trained with supervised objectives since they can exploit different confounders and shortcuts when learning to classify the diverse set of classes.
Second, the distributions ${\bm{R}}({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}})$, ${\bm{R}}({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.35,0.87,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.35,0.87,1}\pgfsys@color@cmyk@stroke{0.65}{0.13}{0}{0}\pgfsys@color@cmyk@fill{0.65}{0.13}{0}{0}\Phi_{\text{Places365}}})$, and ${\bm{R}}({\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}},{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}})$ exhibit similar distribution spreads. Additionally, ${\bm{R}}({\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}},{\color[rgb]{0,0.6,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.6,0}\pgfsys@color@cmyk@stroke{0.64}{0}{0.95}{0.40}\pgfsys@color@cmyk@fill{0.64}{0}{0.95}{0.40}\Phi_{\text{all}}})$ shows a distribution similar to ${\bm{R}}({\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}},{\color[rgb]{0.4,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0,0}\pgfsys@color@cmyk@stroke{0}{0.81}{1}{0.60}\pgfsys@color@cmyk@fill{0}{0.81}{1}{0.60}\Phi_{\text{IN1k}}})$ in Fig. 6. This suggests that including domain-specific models alongside general-purpose ones might not fundamentally alter the representational similarity consistency patterns observed in our analysis.
Our analysis shows that the training data domain significantly influences representational similarity consistency, though mixing domain-specific and general-purpose models preserves the broad distribution patterns of correlation coefficients observed in ImageNet-1k models alone.